What is the mass of a crate with a net force of 300N and accelerate it by 0.750 m/s²

1
Expert's answer
2021-11-16T09:52:14-0500

According to the second Newton's law, the mass is given as follows:


"m = \dfrac{F}{a}"

where "F = 300N" is the net force, and "a= 0.750m\/s^2" is the acceleration. Thus, obtain:


"m = \dfrac{300N}{0.750m/s^2} = 400kg"

Answer. 400 kg.